Timetable Friday, September 13

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304

09.00 – 09.45 Plenary Talk (Chair: Jean-Baptiste Hiriart-Urruty) Adrian Lewis: Robust regularization in optimization

09.50 – 10.20 Minisymposium (Chair: Helmut Maurer) Karsten Eppler: First and second order shape optimization algorithms using wavelet-based BEM (Chair: Irena Lasiecka) Leszek Gasiński: Optimal control problem of Bolza type for evolution hemivariational inequality with hysteresis (Chair: Werner H. Schmidt) Vadim Azhmyakov: Numerical methods for relaxed optimal control problems using proximal-point approach

10.20 – 10.50 Arnd Rösch: Second-order conditions for parabolic control problems with pointwise mixed control-state constraints Vladimir Veliov: Optimal control of uncertain systems with incomplete information for the disturbances Joseph Gergaud: Homotopy methods for a low thrust orbital transfer problem with maximization of final mass

10.50 – 11.20 Nikolai Osmolovskii: Investigation of quadratic forms for broken extremals Nikolaj Smirnov: Discrete multi programmed stabilization of control systems Sjur Didrik Flåm: Nash Play and Newtonian Mechanics

11.20 – 11.45  Coffee break

11.45 – 12.30 Plenary talk (Chair: Rembert Reemtsen) Florian Jarre: Nonlinear semidefinite optimization

12.30 – 13.30  Closing Session